# On-the-Fly OVD Adaptation with FLAME: Few-shot Localization via Active Marginal-Samples Exploration

### v1

@@ Appendix B Proofs
At $(\theta^{\star},\lambda^{\star})$, complementary slackness gives $\lambda_{i}^{\star}=0$ for every $i\notin S$. Hence (8) reduces to
$\nabla_{\theta}\!\left(\tfrac{1}{2}\|\theta\|^{2}-\sum_{i\in S}\lambda_{i}^{\star}\,y_{i}\,\Phi(\theta;x_{i})\right)\Big|_{\theta=\theta^{\star}}=0,$
and together with feasibility and complementary slackness restricted to $i\in S$ this is exactly the KKT system for the reduced problem (7), proving (1).
@@
Lemma B.2 (Support-determination for soft-margin SVM).
$\min_{w,b,\xi}\quad\frac{1}{2}\|w\|^{2}+C\sum_{i=1}^{n}\xi_{i}\quad\text{s.t.}\quad y_{i}(w^{\top}x_{i}+b)\geq 1-\xi_{i},\quad(i=1,\dots,n)$ ($P$)
  1. 1.
    $\min_{w,b,\xi}\quad\frac{1}{2}\|w\|^{2}+C\sum_{i=1}^{n}\xi_{i}\quad\text{s.t.}\quad y_{i}(w^{\top}x_{i}+b)\geq 1-\xi_{i},\quad(i\in S)$ ($P_{S}$)
  2. 2.
@@
Consequently, ($P$) and ($P_{S}$) have the same optimal solutions $(w,b)$. Retraining the soft-margin SVM after removing all non-support points ($i\notin S$) leaves the classifier $x\mapsto\operatorname{sign}(w^{\top}x+b)$ unchanged.
Proof.
KKT system for (P). The Lagrangian is
$\mathcal{L}(w,b,\xi;\alpha,\beta)=\tfrac{1}{2}\|w\|^{2}+C\sum_{i=1}^{n}\xi_{i}-\sum_{i=1}^{n}\alpha_{i}\big(y_{i}(w^{\top}x_{i}+b)-1+\xi_{i}\big)-\sum_{i=1}^{n}\beta_{i}\xi_{i},$
with $\alpha_{i}\geq 0$ (margin) and $\beta_{i}\geq 0$ (nonnegativity). The KKT conditions are
$\nabla_{w}\mathcal{L}=0\ \Rightarrow\ w=\sum_{i=1}^{n}\alpha_{i}y_{i}x_{i},\qquad\partial_{b}\mathcal{L}=0\ \Rightarrow\ \sum_{i=1}^{n}\alpha_{i}y_{i}=0,$
$\partial_{\xi_{i}}\mathcal{L}=0\ \Rightarrow\ \alpha_{i}+\beta_{i}=C\quad(\forall i),$
$y_{i}(w^{\top}x_{i}+b)\geq 1-\xi_{i},\ \ \xi_{i}\geq 0,\quad\alpha_{i}\geq 0,\ \beta_{i}\geq 0,$
$\alpha_{i}\big(1-\xi_{i}-y_{i}(w^{\top}x_{i}+b)\big)=0,\quad\beta_{i}\xi_{i}=0\quad(\forall i).$
Let $(w^{*},b^{*},\xi^{*};\alpha^{*},\beta^{*})$ satisfy these conditions, and set $S:=\{i:\alpha_{i}^{*}>0\}$.
(1) Restriction to $S$.
Consider ($P_{S}$) with Lagrangian
$\mathcal{L}_{S}(w,b,\xi_{S};\alpha_{S},\beta_{S})=\tfrac{1}{2}\|w\|^{2}+C\sum_{i\in S}\xi_{i}-\sum_{i\in S}\alpha_{i}\big(y_{i}(w^{\top}x_{i}+b)-1+\xi_{i}\big)-\sum_{i\in S}\beta_{i}\xi_{i}.$
Since $\alpha_{i}^{*}=0$ for $i\notin S$, the stationarity of (P) at the starred point reduces to
$w^{*}=\sum_{i\in S}\alpha_{i}^{*}y_{i}x_{i},\qquad\sum_{i\in S}\alpha_{i}^{*}y_{i}=0,\qquad\alpha_{i}^{*}+\beta_{i}^{*}=C\ (i\in S),$
which are exactly the stationarity equations of $P_{S}$. Primal/dual feasibility and complementary slackness for $i\in S$ are inherited verbatim. Hence $(w^{*},b^{*},\{\xi_{i}^{*}\}_{i\in S};\{\alpha_{i}^{*},\beta_{i}^{*}\}_{i\in S})$ is a KKT pair for $P_{S}$.
Moreover, for $i\notin S$, $\alpha_{i}^{*}=0$ and $\alpha_{i}^{*}+\beta_{i}^{*}=C$ imply $\beta_{i}^{*}=C$. Then $\beta_{i}^{*}\xi_{i}^{*}=0$ forces $\xi_{i}^{*}=0$, and primal feasibility gives $y_{i}((w^{*})^{\top}x_{i}+b^{*})\geq 1$ for all $i\notin S$; thus the dropped constraints are strictly satisfied at $(w^{*},b^{*})$.
(2) Extension from $S$ to all indices.
Let $(\tilde{w},\tilde{b},\{\tilde{\xi}_{i}\}_{i\in S};\{\tilde{\alpha}_{i},\tilde{\beta}_{i}\}_{i\in S})$ satisfy the KKT system of ($P_{S}$), and define for $i\notin S$:
$\tilde{\alpha}_{i}:=0,\qquad\tilde{\beta}_{i}:=C,\qquad\tilde{\xi}_{i}:=0.$
Then
$\tilde{w}=\sum_{i\in S}\tilde{\alpha}_{i}y_{i}x_{i}=\sum_{i=1}^{n}\tilde{\alpha}_{i}y_{i}x_{i},\qquad\sum_{i=1}^{n}\tilde{\alpha}_{i}y_{i}=\sum_{i\in S}\tilde{\alpha}_{i}y_{i}=0,\qquad\tilde{\alpha}_{i}+\tilde{\beta}_{i}=C\ \ (\forall i),$
so full stationarity holds; dual feasibility is clear; complementary slackness holds on $S$ by assumption and off $S$ because $\tilde{\alpha}_{i}=0$ and $\tilde{\beta}_{i}\tilde{\xi}_{i}=C\cdot 0=0$. If furthermore (as at any full optimum) the removed constraints satisfy $y_{i}(\tilde{w}^{\top}x_{i}+\tilde{b})\geq 1$ for $i\notin S$, then primal feasibility holds for (P) and the extended tuple is a full KKT pair.
Equality of optima and unchanged classifier.
Let $v_{P}$ and $v_{S}$ be the optimal values of (P) and ($P_{S}$). By (1), $(w^{*},b^{*},\{\xi_{i}^{*}\}_{i\in S})$ is feasible for ($P_{S}$) with value
$\tfrac{1}{2}\|w^{*}\|^{2}+C\!\sum_{i\in S}\xi_{i}^{*}\;=\;\tfrac{1}{2}\|w^{*}\|^{2}+C\!\sum_{i=1}^{n}\xi_{i}^{*}\;=\;v_{P},$
so $v_{S}\leq v_{P}$. Conversely, any feasible $(w,b,\{\xi_{i}\}_{i\in S})$ of ($P_{S}$) can be augmented to a feasible point of (P) by setting, for $i\notin S$, $\xi_{i}^{\uparrow}:=\max\{0,\,1-y_{i}(w^{\top}x_{i}+b)\}$, which yields objective $\geq\tfrac{1}{2}\|w\|^{2}+C\sum_{i\in S}\xi_{i}$; hence $v_{P}\leq v_{S}$. Therefore $v_{P}=v_{S}$.
Because the primal objective is strictly convex in $w$, both problems share the same optimal $w$ (unique), and one may choose a consistent optimal bias $b$ (e.g., via any support with $0<\alpha_{i}<C$). Hence removing non-support points and retraining on $S$ leaves $x\mapsto\operatorname{sign}(w^{\top}x+b)$ unchanged. $\Box$
Appendix C Additional experiments
Following, Table 2 provides a detailed per-class breakdown of the Average Precision (AP) on both the DIOR and DOTA datasets, comparing our few-shot method against the zero-shot baseline using the Zero-shot OWL-ViT-v2 fine-tuned on RS-WebLI (appear in second line of Table 1). The missing values in the ’Few-Shot’ columns indicate instances where the initial zero-shot retrieval step failed to find any relevant image embeddings. Without these initial candidates, the few-shot selection process could not proceed, resulting in a method failure for those specific classes."The results highlight the substantial performance gains achieved through our approach across a wide range of object categories.
DIOR Dataset Class Few Shot Zero Shot expressway service area 0.82 0.03 expressway toll station 0.99 0 airplane 0.99 0.84 airport 0 baseball field 0.93 0.62 basketball court 0.87 0.66 bridge 0.49 0.21 chimney 0.94 0.11 dam 0.71 0.04 golf field 0.72 0.01 ground track field 0.79 0.5 harbor 0.64 0.33 overpass 0.75 0.1 ship 0.93 0.72 stadium 0.86 0.57 storage tank 0.68 0.73 tennis court 0.57 0.8 train station 0.01 vehicle 0.79 0.25 windmill 1 0.67
DOTA Dataset Class Few Shot Zero Shot Baseball Diamond 0.88 0.32 Basketball Court 0.83 0.56 Bridge 0.28 0.09 Container Crane 0.95 0.03 Ground Track Field 0.68 0.4 Harbor 0.82 0.36 Helicopter 0.73 0.39 Large Vehicle 0.87 0.32 Plane 0.54 0.78 Roundabout 0.91 0.24 Ship 0.82 0.71 Small Vehicle 0.77 0.28 Soccer Ball Field 0.77 0.48 Storage Tank 0.55 0.79 Swimming Pool 0.58 0.71 Tennis Court 0.77

### v2

@@ Appendix A Proofs
(1) Full $\Rightarrow$ reduced. Let $(w^{\star},b^{\star},\alpha^{\star})$ be any KKT triple for (P), and set $S=\{i:\,y_{i}(w^{\star\top}x_{i}+b^{\star})=1\}$. By complementary slackness, $\alpha_{i}^{\star}=0$ for every $i\notin S$. Hence stationarity reduces to
and together with feasibility and slackness on $S$ these are exactly the KKT conditions of the reduced problem ($\mathrm{P}_{S}$). Thus $\bigl(w^{\star},b^{\star},(\alpha_{i}^{\star})_{i\in S}\bigr)$ is KKT for ($\mathrm{P}_{S}$).
@@
Parts (1)–(2) imply that (P) and ($\mathrm{P}_{S}$) have the same optimal solutions. In particular, removing non–support points leaves the classifier $x\mapsto\operatorname{sign}(w^{\top}x+b)$ unchanged. $\Box$
Proof of Lemma 2.2.
Let $(w^{*},b^{*},\xi^{*};\alpha^{*},\beta^{*})$ be a KKT pair of (P), where the Lagrangian is $\mathcal{L}=\tfrac{1}{2}\|w\|^{2}+C\sum_{i}\xi_{i}-\sum_{i}\alpha_{i}\big(y_{i}(w^{\top}x_{i}+b)-1+\xi_{i}\big)-\sum_{i}\beta_{i}\xi_{i}$ with $\alpha_{i},\beta_{i}\geq 0$ and the implicit constraints $\xi_{i}\geq 0$. The KKT conditions read: (i) $w=\sum_{i}\alpha_{i}y_{i}x_{i}$, $\sum_{i}\alpha_{i}y_{i}=0$, and $\alpha_{i}+\beta_{i}=C$; (ii) $y_{i}(w^{\top}x_{i}+b)\geq 1-\xi_{i}$, $\xi_{i}\geq 0$; (iii) $\alpha_{i}(1-\xi_{i}-y_{i}(w^{\top}x_{i}+b))=0$, $\beta_{i}\xi_{i}=0$. Define $S:=\{i:\alpha_{i}^{*}>0\}$. Since $\alpha_{i}^{*}=0$ for $i\notin S$, the stationarity equations at the starred point reduce to $w^{*}=\sum_{i\in S}\alpha_{i}^{*}y_{i}x_{i}$ and $\sum_{i\in S}\alpha_{i}^{*}y_{i}=0$, while $\alpha_{i}^{*}+\beta_{i}^{*}=C$ holds for $i\in S$. Together with primal/dual feasibility and complementary slackness restricted to $i\in S$, this shows that $(w^{*},b^{*},\{\xi_{i}^{*}\}_{i\in S};\{\alpha_{i}^{*},\beta_{i}^{*}\}_{i\in S})$ satisfies the KKT system of the reduced problem ($\mathrm{P}_{S}$). Moreover, for $i\notin S$ we have $\alpha_{i}^{*}=0$ and thus $\beta_{i}^{*}=C$, which by $\beta_{i}^{*}\xi_{i}^{*}=0$ forces $\xi_{i}^{*}=0$ and hence $y_{i}((w^{*})^{\top}x_{i}+b^{*})\geq 1$, i.e., the dropped constraints are strictly satisfied at $(w^{*},b^{*})$. Conversely, take any KKT pair $(\tilde{w},\tilde{b},\{\tilde{\xi}_{i}\}_{i\in S};\{\tilde{\alpha}_{i},\tilde{\beta}_{i}\}_{i\in S})$ for ($\mathrm{P}_{S}$) and extend by setting $\tilde{\alpha}_{i}:=0$, $\tilde{\beta}_{i}:=C$, $\tilde{\xi}_{i}:=0$ for $i\notin S$. Then $\tilde{w}=\sum_{i\in S}\tilde{\alpha}_{i}y_{i}x_{i}=\sum_{i=1}^{n}\tilde{\alpha}_{i}y_{i}x_{i}\,$ and $\sum_{i=1}^{n}\tilde{\alpha}_{i}y_{i}=0$, while $\tilde{\alpha}_{i}+\tilde{\beta}_{i}=C$ and the complementary slackness equalities hold for all $i$; if $y_{i}(\tilde{w}^{\top}x_{i}+\tilde{b})\geq 1$ for $i\notin S$ (as occurs at any optimum of the full problem), the extension is a full KKT pair for (P). Finally, letting $v_{P}$ and $v_{S}$ be the optimal values of (P) and ($\mathrm{P}_{S}$), the restriction above shows $v_{S}\leq v_{P}$, while any feasible $(w,b,\{\xi_{i}\}_{i\in S})$ of ($\mathrm{P}_{S}$) can be augmented by $\xi_{i}^{\uparrow}:=\max\{0,1-y_{i}(w^{\top}x_{i}+b)\}$ for $i\notin S$ to give a feasible point of (P) with no smaller objective, hence $v_{P}\leq v_{S}$. Thus $v_{P}=v_{S}$, and since the objective is strictly convex in $w$, both problems share the same optimal $w$ (and a consistent $b$), so removing non-support points and retraining leaves the classifier $\operatorname{sign}(w^{\top}x+b)$ unchanged. $\Box$
Proof of Lemma 2.3.
Introduce multipliers $\lambda_{i}\geq 0$ for the constraints in (1). The Lagrangian is
$\mathcal{L}(\theta,\lambda)\;=\;\tfrac{1}{2}\|\theta\|^{2}\;-\;\sum_{i=1}^{n}\lambda_{i}\,y_{i}\,\Phi(\theta;x_{i}),$
and the KKT conditions read
(stationarity) $\theta\;-\;\sum_{i=1}^{n}\lambda_{i}\,y_{i}\,\nabla_{\theta}\Phi(\theta;x_{i})\;=\;0,$
(primal feasibility) $y_{i}\,\Phi(\theta;x_{i})\;\geq\;1\quad(\forall i),$
(dual feasibility) $\lambda_{i}\;\geq\;0\quad(\forall i),$
(complementary slackness) $\lambda_{i}\bigl(y_{i}\,\Phi(\theta;x_{i})-1\bigr)\;=\;0\quad(\forall i).$
(1) Full $\Rightarrow$ reduced. Let $(\theta^{\star},\lambda^{\star})$ be a KKT pair for (1) and $S=\{i:\,y_{i}\,\Phi(\theta^{\star};x_{i})=1\}$. By complementary slackness, $\lambda_{i}^{\star}=0$ for every $i\notin S$, so the stationarity condition reduces to
$\theta^{\star}\;-\;\sum_{i\in S}\lambda_{i}^{\star}\,y_{i}\,\nabla_{\theta}\Phi(\theta^{\star};x_{i})\;=\;0.$
Together with primal/dual feasibility and complementary slackness restricted to $i\in S$, these are precisely the KKT conditions of the reduced problem (2). Hence $(\theta^{\star},(\lambda_{i}^{\star})_{i\in S})$ is KKT for (2).
@@
$\tilde{\theta}\;-\;\sum_{i\in S}\mu_{i}\,y_{i}\,\nabla_{\theta}\Phi(\tilde{\theta};x_{i})\;=\;0,$
[…]
Combining (1)–(2), the KKT solution sets of (1) and (2) coincide. Consequently, removing all non–support points leaves the limiting classifier $x\mapsto\operatorname{sign}(\Phi(\theta;x))$ unchanged. $\Box$